OpenStudy (jerudavis):

WILL MEDAL @REBECCA1588 Variable y varies directly with variable x, and y = 4 when x = 7. What is the value of x when y = 12? _______

OpenStudy (jerudavis):

@Gabylovesyou @Secret-Ninja @Bookworm14 @YoungStudier @Wendy.Ivette11714 @lillylilly1126 @Lydia4goals @mathmale @goldengoodgirl

OpenStudy (jerudavis):

@KendrickLamar2014 @Kidneedshelp

OpenStudy (wendy.ivette11714):

What do you think it is?

OpenStudy (jerudavis):

I REALLY DONT KNO HOLD ON

OpenStudy (wendy.ivette11714):

Ok

OpenStudy (jerudavis):

336 I THINK

OpenStudy (jerudavis):

@superman38

OpenStudy (jerudavis):

@ParthKohli

OpenStudy (wendy.ivette11714):

@KendrickLamar2014

OpenStudy (jerudavis):

U DONT KNO

OpenStudy (wendy.ivette11714):

:O no. I'm not that good with math sorry

OpenStudy (wendy.ivette11714):

@Elsa213

Elsa213 (elsa213):

XD

OpenStudy (wendy.ivette11714):

:P

OpenStudy (jerudavis):

can somebody answer

OpenStudy (jerudavis):

@edwen120

OpenStudy (jerudavis):

@winsteria

OpenStudy (wendy.ivette11714):

@els

OpenStudy (wendy.ivette11714):

@Elsa213

OpenStudy (jerudavis):

@Why_Study_Alone @SamsungFanBoy @Qwertty123 @Askonna_meow @gabbyalicorn @TrojanPoem

OpenStudy (jerudavis):

@TrojanPoem

OpenStudy (trojanpoem):

\[y \alpha x\] \[y = const.x\] \[\frac{ y_{1} }{ y_{2} } = \frac{ x_{1} }{ x_{2} }\] \[\frac{ 4 }{ 12 } = \frac{ 7 }{ x_{2} }\] \[x_{2} = \frac{ 12 . 7 }{ 4 }\] x = 21

OpenStudy (jerudavis):

21

OpenStudy (trojanpoem):

You can also do it directly y got 3 times and it varies with x so x = 3x1

OpenStudy (trojanpoem):

Yep.

OpenStudy (jerudavis):

thanks

OpenStudy (trojanpoem):

Any time.